# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 3.4 Refined invariants of genus two
For genus $g=2$, we restrict our consideration to the case ${\operatorname{Tor}}(P_{\Delta})\simeq{\mathbb{P}}^{2}$. That is, $\Delta$ consists of positive multiples of the vectors $(1,1),(-1,0),(0,-1)$ and defines plane curves of an even degree $2d$. The boundary of the positive quadrant $({\mathbb{P}}^{2}_{\mathbb{R}})^{+}$ consists of the three segments $D_{s}^{+}$, $s=0,1,2$, where $D_{s}^{+}=D_{s}\cap{\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, $D_{s}=\{x_{s}=0\}\subset{\mathbb{P}}^{2}$. We write ${\boldsymbol{w}}_{\partial}={\boldsymbol{w}}_{\partial}^{(1)}\cup{\boldsymbol{w}}_{\partial}^{(2)}\cup{\boldsymbol{w}}_{\partial}^{(0)}$, where
${\boldsymbol{w}}_{\partial}^{(1)}=(w_{i}^{(1)})_{1\leq i\leq n_{1}}\subset D_{1}^{+},\quad{\boldsymbol{w}}_{\partial}^{(2)}=(w_{i}^{(2)})_{1\leq i\leq n_{2}}\subset D_{2}^{+},\quad{\boldsymbol{w}}_{\partial}^{(0)}=(w_{i}^{(0)})_{1\leq i<n_{0}}\subset D_{0}^{+},$
@@
and introduce the completions $\widehat{\boldsymbol{w}}_{\partial}={\boldsymbol{w}}_{\partial}\cup\{w^{(0)}_{n^{0}}\}\in\operatorname{Men}(\Delta)$ and $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$, where ${\boldsymbol{w}}_{in}=\{w_{1},w_{2}\}$,
$w_{1}\in({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}=\left\{\frac{x_{1}}{x_{0}}<0,\ \frac{x_{2}}{x_{0}}>0\right\},\ w_{2}\in({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}=\left\{\frac{x_{1}}{x_{0}}>0,\ \frac{x_{2}}{x_{0}}<0\right\}.$
@@
Denote by $\operatorname{Men}^{\prime}_{2}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$ such that the set ${\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
@@
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{2}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}$
@@
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$. Then we introduce
$\operatorname{Men}_{2}(\Delta)=\left\{{\boldsymbol{w}}\in\operatorname{Men}^{\prime}_{2}(\Delta)\ :\ \overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})\right\}.$
@@
Denote by ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the set of real stable maps of curves of genus $2$ considered up to equivariant isomorphisms. Due to the choice of the points of $\widehat{\boldsymbol{w}}$, each element $[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$, where $S_{2}$ stands for a real Riemann surface of genus $2$, defines an $M$-curve, i.e., ${\mathbb{R}}S_{2}$ consists of three ovals that are mapped to the quadrants ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, respectively. The complement $S_{2}\setminus{\mathbb{R}}S_{2}$ consists of two components, and the choice of one of them, $S_{2}^{+}$ defined a complex orientation on ${\mathbb{R}}S_{2}$ and on ${\Phi}({\mathbb{R}}S_{2})$. Denote by $\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the set of elements of ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,{\boldsymbol{w}})$ equipped with a complex orientation. Similarly to formula (9), we define for each element $\xi=[{\Phi}:(S_{2},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$,
$W_{2}(\xi)=(-1)^{e_{+}(C)}\cdot(-1)^{h_{-}(C)}\cdot\prod_{\begin{array}[]{c}\scriptstyle{k_{i}^{(s)}\equiv 0\mod 2}\\ \scriptstyle{1\leq i\leq n_{s},\ s=0,1,2}\end{array}}{\varepsilon}(\xi,w_{i}^{(s)}),$
@@
where $C={\Phi}(S_{2})$, $e_{+}(C)$ is the number of elliptic nodes which appear in real local equigeneric nodal deformations of singular points of $C$ in $({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, $h_{-}(C)$ is the number of hyperbolic nodes which appear in real local equigeneric nodal deformations of singular points of $C$ in $({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\cup({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$, and ${\varepsilon}(\xi,w_{i}^{(s)})$ equals $1$ or $-1$ according as the complex orientation of the germ ${\Phi}(S_{2},p_{i}^{(s)})$ agrees with the positive orientation of $D_{s}^{+}$ or not. Recall that by [13, Lemma 2.2], the numbers $e_{+}(C)$ and $h_{-}(C)$ are well defined modulo $2$.
@@
For any toric degree $\Delta$ as above and any $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq 2d^{2}$, the number
$W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{2}(\xi)$
does not depend on the choice of a generic sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\mp}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{\pm}$.
We prove this theorem in Section 6.
@@
Note that $\operatorname{Men}_{0}(\Delta)$ is an open dense semialgebraic subset of $\operatorname{Men}^{\tau}_{\mathbb{R}}(\Delta)$. We call its connected components the chambers. The complement to this subset is the union of finitely many semialgebraic strata of codimension $\geq 1$. We call the strata of codimension $1$ the walls.
Now we pick two sequences $\widehat{\boldsymbol{w}}(0)$ and $\widehat{\boldsymbol{w}}(1)$ satisfying the hypotheses of Theorem 3.2. We can join these sequences by a generic path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}\subset\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ which avoids the starta of codimension $\geq 2$. To prove Theorem 3.2, we need to verify the constancy of $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside any given chamber and in all possible wall-crossing events, for all $\kappa$.
@@ 4.1 Invariance inside chambers
Here, we verify that $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ does not change along intervals $t^{\prime}<t<t^{\prime\prime}$ lying in a given chamber. To this end, we show that the projection of ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ to the interval $(t^{\prime},t^{\prime\prime})$ does not have critical points. In such a case, the family ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ is the union of intervals trivially covering $(t^{\prime},t^{\prime\prime})$, and hence it lifts to a trivial family of complex oriented curves and the quantum index persists along each of the components of the latter family.
It is convenient to look at the elements $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ as follows: pick some pair $(\sigma_{0},i_{0})$, where $\sigma_{0}\in P^{1}_{\Delta}$ and $1\leq i_{0}\leq n^{\sigma_{0}}$, and think of $w^{\sigma_{0}}_{i_{0}}$ as a mobile point, whose position is determined by the other points $w^{\sigma}_{i}\in\widehat{\boldsymbol{w}}$ via the Menelaus condition. Put $C={\Phi}_{*}({\mathbb{P}}^{1})$. Observe that the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma)$ at $w^{\sigma}_{i}$ (where $(\sigma,i)\neq(\sigma_{0},i_{0})$) with multiplicity at least $2k^{\sigma}_{i}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma}_{i}}\geq 2k^{\sigma}_{i}\}$. In turn, the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma_{0})$ at a point close to $w^{\sigma_{0}}_{i_{0}}$ with multiplicity at least $2k^{\sigma_{0}}_{i_{0}}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma_{0}}_{i_{0}}}\geq 2k^{\sigma}_{i}-1\}$. Then, the required transversality amounts to the following statement.
@@ 4.2 Walls in the space of constraints
Here, we characterize elements $\xi\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$ that are generic elements of strata of dimension $n-1$ (i.e., of codimension $1$) in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$.
@@
(2) If $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$, where $\widehat{\boldsymbol{w}}(t^{*})$ is a generic element in an $(n-1)$-dimensional stratum in $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$, then $\xi$ is of one of the following types:
@@
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}\simeq{\mathbb{P}}^{1}$, the map ${\Phi}$ is birational onto its image and satisfies the following: either it is smooth at $\widehat{\boldsymbol{w}}(t^{*})$, but has singular branches in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ or it is an immersion everywhere but at one point $w_{i}^{\sigma}(t^{*})$, where it has a singularity of type $A_{2m}$, $m\geq k_{i}^{\sigma}$, and, furthermore, $C={\Phi}(\widehat{C})$ is smooth at each point of $\widehat{\boldsymbol{w}}(t^{*})\setminus\{w_{i}^{\sigma}(t^{*})\}$;
@@
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}$ is one point $p$, each map ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$, $i=1,2$, is either an immersion smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering; furthermore, either ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$ and then all intersection points of the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ are ordinary nodes, or $w={\Phi}(p)\in{\boldsymbol{w}}(t^{*})$, and in the latter case at least one of the maps ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$ is birational onto its image, the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect ${\operatorname{Tor}}(\sigma)$ at $w$ with even multiplicities, and they do not have common point in $\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$;
@@
Proof. (1i) A toric divisor ${\operatorname{Tor}}(\sigma)$ cannot split off alone, since, otherwise, in the deformation along the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$, the intersection points with the neighboring toric divisors would yield points on these toric divisors on the distance less than $\tau$ from the corners of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$. For the same reason we obtain that only all toric divisors together may split off, while their intersection points must smooth out in the deformation. However, this contradicts the rationality of the considered curves.
(1ii) First, note that ${\Phi}(\widehat{C})$ intersects ${\operatorname{Tor}}(\partial P_{\Delta})$ only at $\widehat{\boldsymbol{w}}(t^{*})$. Furthermore, if the images of two irreducible components $\widehat{C}_{1},\widehat{C}_{2}$ of $\widehat{C}$ contain the same point $w_{i}^{\sigma}(t^{*})$ that is different from any other point $w_{j}^{\sigma}(t^{*})$, $j\neq i$, then $w_{i}^{\sigma}(t^{*})={\Phi}(\widehat{C}_{1}\cap\widehat{C}_{2})$ (cf. Lemma 2.4). Due to the genus restriction, it follows that the sequence $\widehat{\boldsymbol{w}}(t^{*})$ satisfies at least three Menelaus conditions, which cuts off $\operatorname{Men}_{\mathbb{R}}^{\tau}(\Delta)$ a polytope of dimension $\leq n-2$, contrary to the dimension $n-1$ assumption.

### v2

@@ 3.4 Refined invariants of genus two
For genus $g=2$, we restrict our attention to the case ${\operatorname{Tor}}(P_{\Delta})\simeq{\mathbb{P}}^{2}$. That is, $\Delta$ consists of positive multiples of the vectors $(1,1),(-1,0),(0,-1)$ and defines plane curves of an even degree $2d$. The boundary of the positive quadrant $({\mathbb{P}}^{2}_{\mathbb{R}})^{+}$ consists of the three segments $D_{s}^{+}$, $s=0,1,2$, where $D_{s}^{+}=D_{s}\cap{\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, $D_{s}=\{x_{s}=0\}\subset{\mathbb{P}}^{2}$ corresponding the sides
$[(0,d),(d,0)],\ [(0,0),(0,d)],\ [(0,0),(d,0)]\in P^{1}_{\Delta},$
@@
respectively. Introduce the completions $\widehat{\boldsymbol{w}}_{\partial}={\boldsymbol{w}}_{\partial}\cup\{w^{(0)}_{n^{0}}\}\in\operatorname{Men}(\Delta)$ and $\widehat{\boldsymbol{w}}=\widehat{\boldsymbol{w}}_{\partial}\cup{\boldsymbol{w}}_{in}$, where ${\boldsymbol{w}}_{in}=\{w_{1},w_{2}\}$,
[…]
@@
Denote by $\operatorname{Men}_{2}(\Delta)$ the set of sequences $\widehat{\boldsymbol{w}}\in\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$ such that the set ${\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ satisfies the conclusions of Lemma 2.4. The closure
@@
$\overline{\bigcup_{\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)}{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}\subset\overline{\mathcal{M}}^{\mathbb{R}}_{2,n}({\operatorname{Tor}}(P_{\Delta}),{\mathcal{L}}_{P_{\Delta}})$
@@
is naturally fibered over the space of sequences $\widehat{\boldsymbol{w}}\subset\operatorname{Men}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$, and we denote the fibers by ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$.
@@
Denote by ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the space of equivariant isomorphism classes of real stable maps of curves of genus $2$. Due to the choice of the points of $\widehat{\boldsymbol{w}}$, each element $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$, where $\widehat{C}$ stands for a real Riemann surface of genus $2$, and $\xi$ defines an $M$-curve, i.e., ${\mathbb{R}}\widehat{C}$ consists of three ovals that are mapped to the quadrants ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}$, ${\operatorname{Cl}}({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$, respectively. The complement $\widehat{C}\setminus{\mathbb{R}}\widehat{C}$ consists of two components, and the choice of one of them, $\widehat{C}^{+}$, defines a complex orientation on ${\mathbb{R}}\widehat{C}$ and on ${\Phi}({\mathbb{R}}\widehat{C})$. Denote by $\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$ the set of elements of ${\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,{\boldsymbol{w}})$ equipped with a complex orientation. Similarly to formula (11), we define for each element $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$,
$W_{2}(\xi)=(-1)^{e_{+}(C)}\cdot(-1)^{h_{-}(C)}\cdot\prod_{\sigma\in P^{1}_{\Delta}}\prod_{k_{i}^{\sigma}\equiv 0\mod 2}{\operatorname{sign}}(\xi,w_{i}^{\sigma}),$
@@
where $C={\Phi}(\widehat{C})$ and $h_{-}(C)$ is the number of hyperbolic nodes which appear in real local equigeneric nodal deformations of singular points of $C$ in $({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}\cup({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$. Put
$W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}W_{2}(\xi),\quad\widetilde{W}^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})=\sum_{\begin{array}[]{c}\scriptstyle{\xi\in\overrightarrow{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})}\\ \scriptstyle{{\operatorname{QI}}(\xi)=\kappa}\end{array}}\widetilde{W}_{2}(\xi),$
where (cf. formula (10)
[…]
@@
Suppose that $n\geq 5$. For any toric degree $\Delta$ as above and any $\kappa\in\frac{1}{2}{\mathbb{Z}}$ such that $|\kappa|\leq 2d^{2}$, the values $W^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}^{\kappa}_{2}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$.
We prove this theorem in Section 6. Since the values $W_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ and $\widetilde{W}_{2}^{\kappa}(\Delta,\widehat{\boldsymbol{w}})$ do not depend on the choice of a sequence $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{2}(\Delta)\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(-,+)}\times({\mathbb{P}}_{\mathbb{R}}^{2})^{(+,-)}$, we use for these values the notations $W_{2}^{\kappa}(\Delta)$ and $\widetilde{W}_{2}^{\kappa}(\Delta)$, respectively.
Remark 3.9
We expect that the statement of Theorem 3.8 holds for the remaining value $n=4$ too. It corresponds to #$\#\Delta=3$ and $T_{\Delta}$ a lattice triangle, while the counted curves of genus $2$ in ${\operatorname{Tor}}(P_{\Delta})$ intersect each toric divisor at one point, where they are unibranch.
@@
Note that $\operatorname{Men}_{0}(\Delta)$ is an open dense semialgebraic subset of $\operatorname{Men}^{\rho}_{\mathbb{R}}(\Delta)$. We call its connected components the chambers. The complement to this subset is the union of finitely many semialgebraic strata of codimension $\geq 1$. We call the strata of codimension $1$ the walls.
Now we pick two sequences $\widehat{\boldsymbol{w}}(0)$ and $\widehat{\boldsymbol{w}}(1)$ satisfying the hypotheses of Theorem 3.2. We can join these sequences by a generic path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}\subset\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ which avoids the loci of codimension $>1$. To prove Theorem 3.2, we need to verify the constancy of $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ and $\widetilde{W}_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ inside any given chamber and in all possible wall-crossing events, for all $\kappa$.
@@ 4.1 Invariance inside chambers
Here, we verify that $W_{0}^{\kappa}(\Delta,\widehat{\boldsymbol{w}}(t))$ does not change along intervals $t^{\prime}<t<t^{\prime\prime}$ for which $\widehat{\boldsymbol{w}}(t)$ lies in a given chamber. To this end, we show that the projection of ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ to the interval $(t^{\prime},t^{\prime\prime})$ does not have critical points. In such a case, the family ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ is the union of intervals trivially covering $(t^{\prime},t^{\prime\prime})$, and hence it lifts to a trivial family of complex oriented curves and the quantum index persists along each of the components of the latter family.
It is convenient to look at the elements $\widehat{\boldsymbol{w}}\in\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ as follows: pick some pair $(\sigma_{0},i_{0})$, where $\sigma_{0}\in P^{1}_{\Delta}$ and $1\leq i_{0}\leq n^{\sigma_{0}}$, and think of $w^{\sigma_{0}}_{i_{0}}$ as a mobile point, whose position is determined by the other points $w^{\sigma}_{i}\in\widehat{\boldsymbol{w}}$ via the Menelaus condition. Put $C={\Phi}_{*}({\mathbb{P}}^{1})$. Observe that the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma)$ at $w^{\sigma}_{i}$ (where $(\sigma,i)\neq(\sigma_{0},i_{0})$) with multiplicity at least $2k^{\sigma}_{i}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma}_{i}}\geq 2k^{\sigma}_{i}\}$. In turn, the tangent space at $C$ to the family of curves $C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|$ intersecting ${\operatorname{Tor}}(\sigma_{0})$ at a point close to $w^{\sigma_{0}}_{i_{0}}$ with multiplicity at least $2k^{\sigma_{0}}_{i_{0}}$ can be identified with the linear system $\{C^{\prime}\in|{\mathcal{L}}_{P_{\Delta}}|\ :\ (C^{\prime}\cdot C)_{w^{\sigma_{0}}_{i_{0}}}\geq 2k^{\sigma}_{i}-1\}$. To ensure the behavior of ${\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{t^{\prime}<t<t^{\prime\prime}}$ as indicated in the previous paragraph, we have to show the transversality of intersection of the tangent space at $C$ to the family of rational curves and the tangent spaces to families of curves matching the above tangency conditions. The required transversality is reformulated in the cohomology language (and proved) in the following lemma.
@@ 4.2 Walls in the space of constraints
Here, we characterize elements $\xi\in\overline{\mathcal{M}}_{0,n}^{{\mathbb{R}},+}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$ such that $\widehat{\boldsymbol{w}}(t)$ is a generic member of a stratum of dimension $n-1$ in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$ (i.e., of codimension $1$ in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$). In view of the results of Section 4.1, we can ignore the walls along which ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{0,n}(\Delta,\widehat{\boldsymbol{w}})$. The following lemma classifies other walls.
@@
(2) If $\xi=[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\operatorname{Tor}}(P_{\Delta})]\in\overline{\mathcal{M}}_{0,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$, where $\widehat{\boldsymbol{w}}(t^{*})$ is a generic element in an $(n-1)$-dimensional stratum in $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)\setminus\operatorname{Men}_{0}(\Delta)$, then $\xi$ is of one of the following types:
@@
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}\simeq{\mathbb{P}}^{1}$, the map ${\Phi}$ is birational onto its image and satisfies the following: either it is smooth at $\widehat{\boldsymbol{w}}(t^{*})$, but has singular branches in ${\operatorname{Tor}}(P_{\Delta})^{\times}$, or $C$ is immersed in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ and at some points $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}(t^{*})$ it has singularity of type $A_{2m}$, $m\geq k_{i}^{\sigma}$, whereas $C$ is smooth at the remaining points of $\widehat{\boldsymbol{w}}(t^{*})$;
@@
    $\widehat{\boldsymbol{w}}(t^{*})$ consists of $n+1$ distinct points, $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $\widehat{C}_{1}\simeq\widehat{C}_{2}\simeq{\mathbb{P}}^{1}$ and $\widehat{C}_{1}\cap\widehat{C}_{2}$ is one point $p$, each map ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$, $i=1,2$, is either an immersion smooth along ${\operatorname{Tor}}(\partial P_{\Delta})$, or a multiple covering of a line intersecting only two toric divisors, while these divisors correspond to opposite parallel sides of $P_{\Delta}$ and the intersection points with these divisors are ramification points of the covering with the maximal ramification index; furthermore, either ${\Phi}(p)\in{\operatorname{Tor}}(P_{\Delta})^{\times}$ and then all intersection points of the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ are ordinary nodes, or $w={\Phi}(p)\in{\boldsymbol{w}}(t^{*})$, and in the latter case at least one of the maps ${\Phi}:\widehat{C}_{i}\to{\operatorname{Tor}}(P_{\Delta})$ is birational onto its image, the curves $C_{1}={\Phi}(\widehat{C}_{1})$, $C_{2}={\Phi}(\widehat{C}_{2})$ intersect ${\operatorname{Tor}}(\sigma)$ at $w$ with even multiplicities, and they do not have common point in $\widehat{\boldsymbol{w}}(t^{*})\setminus\{w\}$;
@@
Proof. (1i) A toric divisor ${\operatorname{Tor}}(\sigma)$ cannot split off alone, since, otherwise, in the deformation along the path $\{\widehat{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$, the intersection points with the neighboring toric divisors would yield points on these toric divisors on the distance less than $\rho$ from the corners of ${\operatorname{Tor}}_{\mathbb{R}}^{+}(P_{\Delta})$. For the same reason we obtain that only all toric divisors together may split off, while their intersection points must smooth out in the deformation. However, this contradicts the rationality of the considered curves.
(1ii) First, note that ${\Phi}(\widehat{C})$ intersects ${\operatorname{Tor}}(\partial P_{\Delta})$ only at $\widehat{\boldsymbol{w}}(t^{*})$. Furthermore, if the images of two irreducible components $\widehat{C}_{1},\widehat{C}_{2}$ of $\widehat{C}$ contain the same point $w_{i}^{\sigma}(t^{*})$ that is different from any other point $w_{j}^{\sigma}(t^{*})$, $j\neq i$, then $w_{i}^{\sigma}(t^{*})={\Phi}(\widehat{C}_{1}\cap\widehat{C}_{2})$ (cf. Lemma 2.4). Due to the genus restriction, it follows that the sequence $\widehat{\boldsymbol{w}}(t^{*})$ satisfies at least three Menelaus conditions, which cuts off $\operatorname{Men}_{\mathbb{R}}^{\rho}(\Delta)$ a polytope of dimension $\leq n-2$, contrary to the dimension $n-1$ assumption.